OpenStudy (anonymous):

Look at this expression. -3x6y4 x-3y-2 Which of the following is the simplest form of this expression?

OpenStudy (anonymous):

of wrt expression

OpenStudy (anonymous):

of what expression

OpenStudy (anonymous):

A. -3x-2y-2 B. -3x2y2 C. -3x9y-2 D. -3x9y6

OpenStudy (anonymous):

ok what is the expression in instructions

OpenStudy (anonymous):

-3x^6y^4/x^-3y^-2

OpenStudy (anonymous):

is the - at the beginning in the midline or right by 3?

OpenStudy (anonymous):

if its in the midline the answer is -3^9y^2

OpenStudy (anonymous):

thx it was correct

OpenStudy (anonymous):

ur welcome

OpenStudy (mathstudent55):

If this is your question, \( \dfrac{-3x^6y^4}{x^{-3}y^{-2}} \) The solution is: \(=-3 x^{6 - (-3)} y^{4 - (-2)} \) \(=-3 x^{6 +3} y^{4 + 2} \) \(=-3 x^9 y^6 \)